OpenStudy (anonymous):

Graph the functions and describe the domain and range of each function. f(x)=-square root x+1

OpenStudy (helder_edwin):

is it \[ f(x)=-\sqrt{x+1} \] ?

OpenStudy (anonymous):

yes!

OpenStudy (helder_edwin):

Ok for a square root to have sense, what is inside has to be (i) positive or zero, or (ii) negative

OpenStudy (anonymous):

ok..

OpenStudy (helder_edwin):

it was a question! choose (i) or (ii)

OpenStudy (anonymous):

positive

OpenStudy (helder_edwin):

positive or zero. remember \[ \sqrt{0}=0 \] Then you have to solve the inequality \[ x+1\geq0 \]

OpenStudy (anonymous):

0?

OpenStudy (helder_edwin):

it that zero or Ohhh?

OpenStudy (anonymous):

zero

OpenStudy (helder_edwin):

it is an inequality! \[ x+1\geq0\qquad\Rightarrow\qquad x\geq-1 \] right?

OpenStudy (anonymous):

yes..

OpenStudy (helder_edwin):

so what is the domain of the function?

OpenStudy (anonymous):

1?

OpenStudy (helder_edwin):

the expression x>=-1 corresponds to an interval what interval?

OpenStudy (anonymous):

-1?

OpenStudy (helder_edwin):

do you know any of these \[ [a,b]\qquad (a.b)\qquad [a,b)\qquad (a.b]\qquad [a,\infty)\qquad (-\infty,b] \] ?

OpenStudy (anonymous):

i recognize them.

OpenStudy (helder_edwin):

do you know it definitions?

OpenStudy (anonymous):

infinite

OpenStudy (helder_edwin):

OK \[ [a,b]=\{x\in\mathbb{R}:a\leq x\leq b\} \] \[ (a,b)=\{x\in\mathbb{R}:a<x< b\} \] \[ [a,\infty)=\{x\in\mathbb{R}:x\geq a\} \]

OpenStudy (helder_edwin):

have you ever seen this?

OpenStudy (anonymous):

yes..but i do NOT know how to identify it. can you show me how to do this first problem so i can maybe try the other ones..

OpenStudy (helder_edwin):

You have to check the right sides of the sets i gave you (what's after the colon). the expression \[ x\geq-1 \] corresponds to which of the three examples i gave u?

OpenStudy (anonymous):

c

OpenStudy (helder_edwin):

Yes! so the domain of the function is \[ D_f=[-1,\infty) \] right?

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

did you understand ?

OpenStudy (anonymous):

a little?

OpenStudy (helder_edwin):

ok we'll get to it later. now for the range of the function first solve for x the equation \[ y=f(x)=-\sqrt{x+1} \]

OpenStudy (helder_edwin):

post what you get

OpenStudy (anonymous):

WHAT WOULD I PUT FOR X?!

OpenStudy (helder_edwin):

nothing! just x!

OpenStudy (anonymous):

ok..now what.

OpenStudy (helder_edwin):

what did you get?

OpenStudy (anonymous):

if X is nothing then what am i solving for! +1?

OpenStudy (helder_edwin):

let me do it! you just write x! \[ \LARGE y=f(x)=-\sqrt{x+1} \] \[ \LARGE -y=\sqrt{x+1} \] \[ \LARGE y^2=(-y)^2=x+1 \] \[ \LARGE x=y^2-1 \] OK?

OpenStudy (anonymous):

..yeah what is x

OpenStudy (helder_edwin):

just x!

OpenStudy (anonymous):

you keep saying write x..but x isnt anything

OpenStudy (helder_edwin):

x and y are variables, y depends on x, and x is independent right?

OpenStudy (anonymous):

ya

OpenStudy (helder_edwin):

when you solve the equation i solved you are finding the inverse function (or relation if f is not inyective)

OpenStudy (helder_edwin):

now in the expression \[ \LARGE -y=\sqrt{x+1} \] we have a restriction \[ \LARGE \sqrt{\text{?}}\geq0 \] so this means that \[ \LARGE -y\geq0\Rightarrow y\leq0 \] this gives us the range \[ \LARGE R_f=(-\infty,0] \]

OpenStudy (helder_edwin):

are there? tired or bored?

OpenStudy (anonymous):

sorry..i stepped out for a while

OpenStudy (helder_edwin):

no problem did you understand?